OpenStudy (anonymous):

Rizza and Angel are 30km apart. If they leave at the same time and walk in the same direction, Rizza overtakes Angel in 6 hours. If they walk towards each other, they meet in 5 hours. What are their speeds?

OpenStudy (ganpat):

s = d/ t let Rizza overtake Angel at x SO Sr = x/6.. and Sa = (x-30)/6 Walking towards each other.. (x/6) + (x-30) = 30/5 5x + 5x - 150 = 180 10x = 330 x = 33... so they over take at 33.. Speed Rizza = 33/6 = 5.5 kmph and Speed Angela = 3/6 = 0.5 kmph I think so.. not very sure...

OpenStudy (anonymous):

thank you

OpenStudy (ganpat):

is that a answer ??